OpenStudy (anonymous):

lim ((x^2 + 3)-(-5^2 + 3)) / (x--5) x->-5

OpenStudy (jamesj):

I assume you just learnt the definition of derivative? In which case, what does this quotient and limit look like?

OpenStudy (anonymous):

ya so its f(x)=x^2+3 , P(-5,28)

OpenStudy (jamesj):

Right, so this limit is the derivative of f evaluated at -5. So calculate the derivative and then evaluate it at -5

OpenStudy (anonymous):

yes but the problem I'm having is with the algebra and the simplification of the limit

OpenStudy (jamesj):

No, don't use the limit. Just look at this function. What is the derivative with respect to x of x^2 ? And of 25? Hence of f(x)?

OpenStudy (jamesj):

Or does the question ask you explicitly to evaluate the limit?

OpenStudy (jamesj):

For the record, the derivative of f(x), f'(x) = 2x + 0 = 2x. So f'(-5) = 2.-5 = -10.

OpenStudy (anonymous):

well I'm trying to find the slope of the tangent line

OpenStudy (jamesj):

Yes, I know. I just gave you one way to get the answer and the answer itself. But calculating now directly, ((x^2 + 3)-(-5^2 + 3)) / (x--5) = (x^2 -25)/(x+5) = (x+5)(x-5)/(x+5) = x - 5 Hence the limit as x->-5 of ((x^2 + 3)-(-5^2 + 3)) / (x--5) is the limit as x->-5 of (x-5) I.e., -10.

OpenStudy (anonymous):

oh wow thank you